OpenStudy (cutiecomittee123):

a few more trig questions that i need help with sin(x)(sin(x)-1) How do I solve this

OpenStudy (mathstudent55):

What is that? Is it an equation? Where is the equal sign?

OpenStudy (cutiecomittee123):

oh yeah it is equal to 0 it should be sin(x)(sin(x)-1)=0

OpenStudy (mathstudent55):

The algebraic part of this equation is similar to: y(y - 1) = 0 This is a quadratic equation that has already been factored, so just set each factor equal to zero to find y. In my example, you'd continue: y = 0 or y - 1 = 0 y = 0 or y = 1 Do you understand my example?

OpenStudy (mathstudent55):

Your equation is similar to my example in the algebra part, but then you need to continue with the trig part.

OpenStudy (cutiecomittee123):

so just multiply it out

OpenStudy (mathstudent55):

No. Just use the zero product rule: \(\sin x (\sin x-1)=0\) You have a product equaling zero. A product is only equal to zero if one of its factors equal zero. That is the zero product rule. Set each factor equal to zero and solve each one as an equation. \(\sin x = 0\) or \(\sin x - 1 = 0\)

OpenStudy (cutiecomittee123):

i am kinds stuck though sin(x)^2-sin(x)-1=0 is what i get when i do that

OpenStudy (cutiecomittee123):

oh i see it okay

OpenStudy (mathstudent55):

What you did is as if you got a factored quadratic equation, and instead of setting each factor equal to zero, you multiplied it out.

OpenStudy (cutiecomittee123):

okay so sin(x)=0 and sin(x)=1

OpenStudy (mathstudent55):

Some of the work was already done for you. You didn't have to factor it. Now set each factor equal to zero and solve each factor as an equation.

OpenStudy (mathstudent55):

Exactly. Where the sine = 0?

OpenStudy (mathstudent55):



OpenStudy (cutiecomittee123):

at 0 and pi

OpenStudy (mathstudent55):

Great. Now the second equation. \(\sin x = 1\)

OpenStudy (mathstudent55):